I know I need to use substitution to solve the system, but can you help me here.
4x-4y=-8
y=2+x
please solve

Answer by funmath(2933)

Use substitution to solve the system:
4x-4y=-8
y=2+x
Let y=2+x in the first equation:
4x-4(2+x)=-8
4x-8-4x=-8
-8=-8
Because the variable was eliminated and -8=-8 this is a dependent system, the solution is all real numbers because algebraically no matter what number we let x = they will both have the same y. Graphically these are two lines lying on top of each other having all points in common.
Some teachers let you say the solution is ALL REAL NUMBERS
Others prefer coordinate notation (x,y)=(x,2+x)
Happy Calculating!!!